OpenStudy (anonymous):

Suppose a researcher estimates that the initial population of a colony of cells is 100. If the cells reproduce at a rate of 25% per week, what is the expected population of the colony in six weeks? Use the formula: N = N(1 + r)^t

ganeshie8 (ganeshie8):

its similar to the previous one... any ideas what r, N, t are here ?

OpenStudy (anonymous):

r=.25

OpenStudy (anonymous):

t=6

OpenStudy (anonymous):

is N 100?

ganeshie8 (ganeshie8):

good start :)

OpenStudy (anonymous):

which N?

ganeshie8 (ganeshie8):

right side N = 100. formul is wrong it is using same variable for Final & Current population lets call Nf the final population : Nf = N(1 + r)^t N = 100, r = .25 t =6 and find the Final population in 6 weeks

ganeshie8 (ganeshie8):

Nf = 100(1+.25)^6

OpenStudy (anonymous):

THANKS!

ganeshie8 (ganeshie8):

np :)